OpenStudy (anonymous):

are the steps for this equation correct? equation follows: 2 log_2 2 + 2 log_2 6 - log_2 3x = 3 log_2 2^2 + log_2 6^2- log_2 3x = 3 log_2 ((2^2*6^2)/3x)=3 log_2 ((4*36)/3x)=3 log_2 (144/3x)=3 but I get stuck after this. And I'm not even sure if this is correct. someone please help? :)

OpenStudy (fibonaccichick666):

just to check, this is your work? \[2 log_2 2 + 2 log_2 6 - log_2 3x = 3 log_2 2^2 + log_2 6^2- log_2 3x = 3 \]\[log_2 ((2^2*6^2)/3x)=3 log_2 ((4*36)/3x)=3 log_2 (144/3x)=3\]

OpenStudy (anonymous):

yeah thats it :)

OpenStudy (fibonaccichick666):

ok so, why in the second line, is your 2and 6 squared?

OpenStudy (fibonaccichick666):

then why do you have a 3 as a coefficient?

OpenStudy (anonymous):

because i was following the third rule of logarithmic functions http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif

OpenStudy (fibonaccichick666):

good, (I want you to prove your steps to me just like that when I ask these sort of weird questions. I swear there is a point to it)

OpenStudy (anonymous):

because that is how the equation was written

OpenStudy (anonymous):

haha alright

OpenStudy (fibonaccichick666):

? wait, what is the original? i thought it was just \[2 log_2 2 + 2 log_2 6 - log_2 3x = 3 \]

OpenStudy (fibonaccichick666):

oops, nvm saw the 3 from the previous step

OpenStudy (anonymous):

oh haha ok

OpenStudy (fibonaccichick666):

so that first step is correct now can you justify \[2 log_2 2 + 2 log_2 6 - log_2 3x = 3~~~ log_2 2^2 + log_2 6^2- log_2 3x = 3~~~ \color{red}{log_2 ((2^2*6^2)/3x)=3} \]

OpenStudy (fibonaccichick666):

\[2 log_2 2 + 2 log_2 6 - log_2 3x = 3~~~\] \[log_2 2^2 + log_2 6^2- log_2 3x = 3~~~\] \[\color{red}{log_2 ((2^2*6^2)/3x)=3}\]

OpenStudy (fibonaccichick666):

(I trust you can do the basic algebra involved, so just reference the rules and a brief exp)

OpenStudy (anonymous):

right. i got stuck on the fact that i have to solve for x in this equation. i simplified it down to \[\log_{2}\left(\begin{matrix}144/3x\end{matrix}\right)=3\] but i dont know what the next step would be to get x by itself. do i get rid of the log or simplify the parenthesis (which confuses me)

OpenStudy (fibonaccichick666):

well if 144 is whole when you divide by 3 then you can simplify, but we can now get rid of that log

OpenStudy (fibonaccichick666):

so if you had \(e^{ln(x)}=?\) could you simplify that?

OpenStudy (anonymous):

no....lol that looks like complete gibberish. but if i have \[144\div3x=3\] i could multiply both sides by 3x and have \[144=9x\] and simplify it from thee to be \[x=16\] is this correct?

OpenStudy (fibonaccichick666):

no can't do that

OpenStudy (fibonaccichick666):

logs are functions

OpenStudy (fibonaccichick666):

you can't just pull out the parentheses

OpenStudy (fibonaccichick666):

ie it could just be \(log_2 x=3\) solve for x

OpenStudy (fibonaccichick666):

it's a one step solution, so anywho watch this: https://www.khanacademy.org/math/algebra2/logarithms-tutorial/logarithm_properties/v/introduction-to-logarithm-properties

OpenStudy (fibonaccichick666):

then check out the inverse of a log

OpenStudy (anonymous):

but what happens to the 144/3x if you turn it into \[\log_{2} x=3\]????

OpenStudy (fibonaccichick666):

not yet, that was a counterex

OpenStudy (fibonaccichick666):

just an easier ex to get the basic concept accross

OpenStudy (anonymous):

oh!!!! i figured it out! thank you :D

OpenStudy (fibonaccichick666):

k show me your work so i can make sure you get it

OpenStudy (anonymous):

x=6 because you change the log formula to an exponential formula so that it reads\[2^{3}=144\div3x\] then multiply both sides by 3x and you end up with\[24x=144\] then simplify and then you get\[x=6\]

OpenStudy (anonymous):

thats right, isnt it :)

OpenStudy (fibonaccichick666):

uhm, yea, not how I think about it but that works

OpenStudy (fibonaccichick666):

also, if you divide by 3x the x is on the bottom, check your algebra

OpenStudy (anonymous):

oh, right..in fixed it

OpenStudy (fibonaccichick666):

show me again

OpenStudy (anonymous):

\[2^3=144\div3x\] multiply both sides by the inverse of 3x\[(1/3x)*2^3=144\div3x*(1/3x)\] im pretty sure this is correct because then the x wont be on the bottom right?

OpenStudy (fibonaccichick666):

no x is still on the bottom

OpenStudy (anonymous):

argh im confused

OpenStudy (fibonaccichick666):

how about get rid of the numbers and leave the x alone?

OpenStudy (fibonaccichick666):

ie you have 4=6v solve for v what do you do?

OpenStudy (anonymous):

okay...so um\[8=144\div3x\]\[8=(144\div3)x?\]

OpenStudy (fibonaccichick666):

yup

OpenStudy (fibonaccichick666):

now get rid of the blah in front of the x